2. Lets pretend that you came into lab and found a container marked 0.030 M Weak Base. After calibrating a pH electrode, y
Answer #1

Given weak base concentration, C = 0.08M Weak base por = 7 (pky - load o . = 10.50, But we know PH 4 pokalu a pole = 14-pH =